[5 5 5 5 20 pts] A system is defined as shown on the right y[e]=[a1 -1 =3] .[e]=[31-13] 1=0 ן;ria Irrזן וחןa) Find Y[k] & X[k] b) Find H[k](Yes H[K]) c) Find h[n] d) Find y[n]=x[n]®, h[n]
Answer #1

hel yeni = [-2. 1,-1,- 3} 2011 = (3.1,-1,3] y (n=-28(n-u + 8 (0-2) – 8 (n-3) -38(0-4) X(n) = 2 sins + Slues) -sin-21 43 8(n-3H(2) = -27342-7-1 Z(2 2²4 22z+32 - 2²+0.52-052-25 z (2+0.5220.52+1.5) - 23+ 0.53 0.57-1.5 2 (2+1.5) (2 - 2+1) (2) - -2 0.52-0kon) = -2866) * gina-1) + 2 sima@gen-2) -35lm, 8(n-3) S(n-1)=(n-1) + 8(N-1) 48(1-2) – 1.58cm 128 123) + 81n-24* Sin 1 -S(n-23